# Cosmic Cartography II: completing galaxy catalogs for gravitational-wave cosmology

### v1

@@ 2.3 Galaxy bias
Numerical simulations of galaxy formation show that galaxy density does not scale linearly with the DM density contrast. Instead, whether a galaxy is formed is a complex process of gravitational dynamics, baryonic feedback, and hydrodynamic processes [69, 70, 71, 72, 73, 74, 75]. In this work, we adopt a simplified model by assuming that the DM density contrast within each voxel is directly related to the galaxy formation rate via a function, slightly modified from [76], given by
@@ 2.4 Magnitude distribution
At first glance, there seems to be a strong degeneracy between the measurement of the magnitude distribution and the estimation of the missing galaxies. However, as we will now argue, the high completeness of the survey at low redshift can break this degeneracy. Indeed, nearby voxels (low redshift) inform on the absolute magnitude distribution, up to a limit that corresponds to the faintest galaxy that can be observed near-by. As we move to larger redshifts (cf. Fig. 5), the apparent magnitude distribution shifts (following the blue line of Fig. 5) to larger values (i.e. galaxies appear at higher apparent magnitudes for a fixed absolute magnitude). For voxels at large redshift a significant portion of galaxies is too faint to be observed. However, from the inferred magnitude distribution at low redshifts and the abundance of bright galaxies we can infer the number of faint galaxies missed. From this argument, it is clear that the redshift (in)dependence of the magnitude distribution is an important systematic for the reconstruction.
In order to be agnostic, we adopt a flexible magnitude distribution that does not rely on a specific functional form (as opposed to, for example, a Schechter distribution [77]). This flexible approach is based on a Gaussian random field, which necessitates the assumption of a two-point function which determines how strongly and how fast the PDF fluctuates in magnitude space. We detail the implementation of this approach in Appendix A.1 and show examples of possible magnitude distributions in Fig. 18. Let us stress here that this distribution is assumed to be redshift-independent.
@@ 2.5.3 HEALPix binning
We utilize the HEALPix (Hierarchical Equal Area isoLatitude Pixelization) framework [80, 81] to divide the celestial sphere into equal-area pixels. The HEALPix parameter nside controls the number of two-dimensional segments via #$\#\text{pixels}=12\,\texttt{nside}^{2}$, i.e. the map resolution. For our analyses, we choose $\texttt{nside}=16$, which divides the sky in 3072 segments, each with an area of $\sim 13\,{\rm deg}^{2}$. While it is desirable to have a higher resolution, we are limited by the cartesian grid used to generate the DM density contrast. For a simulation box of size $\mathcal{O}(6000)$ Mpc, we require $\mathcal{O}(10^{2})$ grid points per dimension for $\texttt{nside}=16$, which is $\sim 10^{6}$ pixels, i.e. parameters to be estimated. In principle, the inference algorithm could handle larger parameter spaces but the model size is currently limited by the GPU memory. Future multi-GPU inference development will thus allow to improve the sky discretization to higher nside.
@@ 2.6 Redshift uncertainties
Note that we introduce a correction term that counter-acts the effects at the boundary of our redshift range for the numerical evaluation of Eq. (2.11). This correction enforces that a constant $\mathbb{P}(z)$ results in a constant $\mathbb{P}(\hat{z})$. Otherwise, boundary regions would have lower probability density, due to the absence of galaxies outside the simulated redshift domain that would otherwise scatter into it, leading to an apparent deficit at the boundaries.
@@ 2.7 Detection probability
The detection probability of galaxies is a key ingredient in the reconstruction of the full galaxy field. While this is a complicated function of the data (e.g. the image, or spectrum), the classification as “galaxy” or “not galaxy” is typically approximated as a function dependent on the observed fluxes, possibly in multiple bands. We take the detection probability here as a function of the apparent magnitude in one given band and the sky position. Since these two quantities are not enough to determine with certainty whether the galaxy is detected or not, the detection probability becomes a continuous function between 00 and $1$. [^5] While the inclusion of the data point is a binary decision as a function of the data, it is a continuous function as a function of its true parameters, since one marginalizes over different noise realizations.
@@
where “det” stands for “detected” and $\lambda$ is a set of galaxy parameters that is not included in our survey data. While the detection probability could depend on a wider set of galaxy parameters $\lambda$ we assume that the detection probability can be effectively modeled solely as a function of the sky position $\varphi_{\perp}$, and apparent magnitude $m$, i.e. $\mathbb{P}({\rm det}|\varphi_{\perp},m)$.
@@ 2.8 Likelihood
Putting Eq. (2.3), (2.6) and (2.11) together, in combination with the modeling of the detection probability, we can combine the relevant terms of the above sections to compute the rate of detected galaxies with a measured redshift $\hat{z}$, a sky position $\varphi_{\perp}$, and an apparent magnitude $m$ as
@@ 2.10 Sampling
Generating posterior samples from a high-dimensional posterior distribution is a challenging task because the posterior tends to occupy a small volume of the overall prior volume. For this reason we make use of Hamiltonian Monte Carlo (HMC) sampling, which proposes new samples of an MCMC chain by integrating trajectories through the space of parameters, ensuring proposed samples to have high acceptance probability. The integration of paths requires taking gradients, which we can easily compute, as our model is implemented using the auto-differentiable programming package jax [82].
We generate our samples using a particular type of HMC, the No-U-turn sampler (NUTS) [83, 84, 85] that is implemented in the numpyro package [86]. NUTS extends HMC to generate fairly uncorrelated samples by identifying a U-Turn in the trajectories of the HMC sampling process. No burn-in removal is required for this sampler as this accounted for by the so-called warm-up phase. During this phase (for which we choose a length of 2000 samples), one tunes the step-size and mass matrices of the integrator. We then draw 1500 posterior samples of our target parameters with three chains.
@@ 3 Simulated data
To validate our method on realistic data, we use the mock galaxy catalogs of [87], which are based on the Millennium Simulation DM density field [88]. The original catalog represents a cubic volume of $(500\,{\rm Mpc}/h)^{3}$. Since we want to build a comparable catalog to glade+ [63], this coverage is insufficient for our analysis (since glade+ covers a much larger volume), so we replicate the catalog in an $9\times 9\times 9$ tiling (including random rotations), effectively increasing the total volume to $(4500\,{\rm Mpc}/h)^{3}$, while maintaining the large-scale statistical properties of the distribution. For this tiling, we use the same snapshot of the Millennium simulation at $z=0$. To match the overall number of observed galaxies in glade+, we select only $1\%$ of all available galaxies. This yields approximately a total of $\mathcal{O}(3\times 10^{6})$ galaxies. Assuming a $\Lambda$CDM cosmology with $H_{0}=67\,\,{\rm km\,s^{-1}\,Mpc^{-1}}$ and $\Omega_{{\rm m}}=0.3$, we convert the galaxy comoving coordinates to redshift and sky position.
@@ 4.1.1 Validating the reconstruction (redshift + polar angle)
We start with marginalizing over the absolute magnitude coordinate [^7] We reconstruct the number counts up to an absolute magnitude of $M_{\rm thresh}=-19$. We determine in Sec. 4.4 the absolute magnitude above which the reconstruction can no longer be trusted. , and fix the azimuthal angle to 0 and $\pi$ in order to obtain the entire $x-z$ plane, shown in Fig. 7. The panels are, from left to right, top to bottom: (1) The true galaxy number count, galaxies binned by their true redshift and polar angle, (2) the observed galaxy number count, binned by their measured redshift and polar angle, (3) the reconstructed galaxy Poisson rate in true redshift and polar angle, (4-6) the galaxy densities (same as panel (1-3) but dividing out the comoving volume), (7) the relative error of the galaxy Poisson rate, i.e.
[…]
@@
[…]
@@
[…]
@@ 4.1.4 The reconstructed line-of-sight redshift prior
To understand the implications of the reconstruction, we plot the true galaxy number count and inferred galaxy Poisson rate for a given HEALPixel, summing over the absolute magnitude counts. We focus on two different cases: (i) a pixel with a large number of observed galaxies, i.e. with a large apparent magnitude threshold, and (ii) a pixel in the “milky-way region”, i.e. with a small apparent magnitude threshold. The two cases are shown in the top and and bottom panels of Fig. 10, respectively. There are two differences between the true and observed number counts: while they agree for low redshifts, the ratio of observed to true galaxy number counts drops for larger redshift due to fainter galaxies being missed. The second difference is that the observed number counts lack the small-scale fluctuations of the true number counts—this is due to the redshift uncertainties smearing out any fluctuations smaller than the standard deviation of the redshift distribution.
@@
Figure 11: Galaxy number counts in apparent magnitude and redshift for a selected pixel (top) and summed over the full sky (bottom). The columns from left to right are: the true number counts, the observed number count, the reconstructed Poisson rate, the relative difference, i.e. Eq. (4.4). The cut in the observed number counts at $m=22$ is due to the likelihood being evaluated on a grid up to this value – very few galaxies are observed above these apparent magnitudes. Note that the upper-left dark triangle is due to us excluding all magnitudes above $-19$, while the lower-right section is dark because no galaxies above a certain brightness exist. During inference, we do not impose any upper threshold on the absolute magnitude, this is only applied in post-processing.
@@ 4.1.5 The inferred absolute magnitude distribution
Let us first focus on the true catalog (first column). We see the apparent magnitude evolution (the approximately square root trend), reminiscent of Eq. (2.8) and Fig. 5. As galaxies are located at larger redshifts, they appear fainter (i.e. at larger $m$). We also notice an overall increase in the number counts with redshift, which is due to the larger associated comoving volume for voxels at larger redshifts. The second column shows the observed galaxy number counts, which are smaller than the true number counts due to the lower detection probability at larger apparent magnitudes. We have marked the apparent magnitude at which only $50\%$ of the galaxies are detected as a red dashed horizontal line. Above this line only few galaxies are detected – at an apparent magnitude of $m=22$ the observed catalog is essentially empty. Note that we have applied an absolute magnitude cut of $M_{\rm thr}=-19$, which is reflected in a dark region in the top left region of each plot. We decide to eliminate these galaxies (in post-processing) since they have low GW host probability and hence also contribute little to the redshift measurement for GWs. We will explore the impact of this cut in more detail in Sec. 4.4.
@@
Fig. 12 shows the absolute magnitude distribution of all (detected) galaxies in light blue (green), and the reconstructed galaxy Poisson rate in dark blue. The reconstruction recovers the full distribution faithfully, up to faint galaxies ($M_{h}\sim-18$), where $M_{h}$ is defined in Eq. (A.1). Since we reconstruct the galaxy rate only up to magnitudes above this value, the deviation between the true distribution and our reconstruction does not impact the final estimate of the number counts, as discussed in Sec. 4.4. The faithful recovery of the true magnitude distribution also justifies the generic galaxy magnitude modeling through a correlated Gaussian field that was introduced in App. A.1. It is universal enough to capture the non-trivial galaxy magnitude distribution of [87].
@@ 4.3 Impact of redshift uncertainties
For the simulation of the observed galaxy number counts, we assumed a Gaussian distribution for the redshift uncertainties, with an evolving standard deviation of $\sigma_{z}=0.01\times(1+z)$. This is a reasonable assumption for upcoming spectroscopic surveys [89, 90, 91], but it is important to understand how the results depend on the modeled redshift uncertainty as we assume it is constant. To this end, we repeat the analysis of Sec. 4.1 for different values of $\sigma_{z}$ (fixing $H_{0}=67\,\,{\rm km\,s^{-1}\,Mpc^{-1}}$). We assume (i) $\sigma_{z}=0.005$, (ii) $\sigma_{z}=0.02$ and (iii) $\sigma_{z}=0.07$, using the same simulated redshift uncertainties, i.e. $\sigma_{z}=0.01\times(1+z)$ in all cases. In Fig. 15, we show the deviations $\Delta_{\rm std}$ (cf. Eq. (4.3)) for each of these three analyses. As explained in a previous footnote, the $\Delta_{\rm std}$ distribution should follow approximately a zero-mean, unit-variance normal distribution. In the case of $\sigma_{z}=0.005$ (violet), the variance of the deviations is too large, indicating that the galaxy rate uncertainties are underestimated, as expected: the redshift uncertainties are too small when compared to the simulated uncertainties. In the case of $\sigma_{z}=0.02$ (blue), the deviations are significantly closer to the zero-mean unit-variance Gaussian distribution. Finally, in the case of $\sigma_{z}=0.07$ (yellow), the deviations show a variance that is too small – the uncertainties are overestimated.
We thus conclude that in order to be conservative, one should either (i) account for the redshift-dependent uncertainties or, (ii) use a larger redshift uncertainty overall that the effectively accounts for the redshift-evolving uncertainties. While the first option is the most accurate, it is also the most computationally expensive as the process to compute $\mu(\hat{z},\varphi_{\perp},m;\Lambda)$ will no longer be a convolution with a Gaussian kernel (cf. Eq. (2.11)).

### v3

@@ 2.3 Galaxy bias
Numerical simulations of galaxy formation show that galaxy density does not scale linearly with the DM density contrast. Instead, whether a galaxy is formed is a complex process of gravitational dynamics, baryonic feedback, and hydrodynamic processes [Kuhlen:2012ft, Reddick:2012qy, Zheng:2015iia, Chaves-Montero:2015iga, Zavala:2019gpq, deMartino:2020gfi, Kokron:2021xgh]. In this work, we adopt a simplified model by assuming that the DM density contrast within each voxel is directly related to the galaxy formation rate via a function, slightly modified from [Neyrinck:2013ezr], given by
@@ 2.4 Magnitude distribution
At first glance, there seems to be a strong degeneracy between the measurement of the magnitude distribution and the estimation of the missing galaxies. However, as we will now argue, the high completeness of the survey at low redshift can break this degeneracy. Indeed, nearby voxels (low redshift) inform on the absolute magnitude distribution, up to a limit that corresponds to the faintest galaxy that can be observed near-by. As we move to larger redshifts (cf. Fig. 5), the apparent magnitude distribution shifts (following the cyan line of Fig. 5) to larger values (i.e. galaxies appear at higher apparent magnitudes for a fixed absolute magnitude). For voxels at large redshift a significant portion of galaxies is too faint to be observed. However, from the inferred magnitude distribution at low redshifts and the abundance of bright galaxies we can infer the number of faint galaxies missed. From this argument, it is clear that the redshift (in)dependence of the magnitude distribution is an important systematic for the reconstruction.
In order to be agnostic, we adopt a flexible magnitude distribution that does not rely on a specific functional form (as opposed to, for example, a Schechter distribution [Schechter:1976iz]). This flexible approach is based on a Gaussian random field, which necessitates the assumption of a two-point function which determines how strongly and how fast the PDF fluctuates in magnitude space. We detail the implementation of this approach in Appendix A.1 and show examples of possible magnitude distributions in Fig. 18. Let us stress here that this distribution is assumed to be redshift-independent and independent from the DM density contrast.
@@ 2.5.3 HEALPix binning
We utilize the HEALPix (Hierarchical Equal Area isoLatitude Pixelization) framework [Gorski:2004by, Zonca:2019vzt] to divide the celestial sphere into equal-area pixels. The HEALPix parameter nside controls the number of two-dimensional segments via $\text{pixels}=12\,\texttt{nside}^{2}$, i.e. the map resolution. For our analyses, we choose $\texttt{nside}=16$, which divides the sky in 3072 segments, each with an area of $\sim 13\,{\rm deg}^{2}$. While it is desirable to have a higher resolution, we are limited by the cartesian grid used to generate the DM density contrast. For a simulation box of size $\mathcal{O}(6000)$ Mpc, we require $\mathcal{O}(10^{2})$ grid points per dimension for $\texttt{nside}=16$, which is $\sim 10^{6}$ pixels, i.e. parameters to be estimated. In principle, the inference algorithm could handle larger parameter spaces but the model size is currently limited by the GPU memory. Future multi-GPU inference development will thus allow to improve the sky discretization to higher nside.
@@ 2.6 Redshift uncertainties
[…]
@@ 2.7 Detection probability
The detection probability of galaxies is a key ingredient in the reconstruction of the full galaxy field. While this is a complicated function of the data (e.g. the image, or spectrum), the classification as “galaxy” or “not galaxy” is typically approximated as a function dependent on the observed fluxes, possibly in multiple bands. We take the detection probability here as a function of the apparent magnitude in one given band and the sky position. Since these two quantities are not enough to determine with certainty whether the galaxy is detected or not, the detection probability becomes a continuous function between 0 and 11.444While the inclusion of the data point is a binary decision of the data, it is a continuous function as a function of its true parameters, since one marginalizes over different noise realizations.
@@
where “det” stands for “detected” and $\lambda$ is a set of galaxy parameters that is not included in our survey data. While the detection probability could depend on a wider set of galaxy parameters $\lambda$ we assume that the detection probability can be effectively modeled solely of the sky position $\varphi_{\perp}$, and apparent magnitude $m$, i.e. $\mathbb{P}({\rm det}|\varphi_{\perp},m)$.
@@ 2.8 Likelihood
Putting Eq. (2.3), (2.6) and (2.11) together, in combination with the modeling of the detection probability, we can combine the relevant terms of the above sections to compute the rate of detected galaxies at a given measured redshift $\hat{z}$, a sky position $\varphi_{\perp}$, and an apparent magnitude $m$ as
@@ 2.10 Sampling
Generating posterior samples from a high-dimensional posterior distribution is a challenging task because the posterior tends to occupy a small volume of the overall prior volume. For this reason we make use of Hamiltonian Monte Carlo (HMC) sampling, which proposes new samples of an MCMC chain by integrating trajectories through the space of parameters, ensuring proposed samples to have high acceptance probability. The integration of paths requires taking gradients, which we can easily compute, as our model is implemented using the auto-differentiable programming package jax [jax2018github].
We generate our samples using a particular type of HMC, the No-U-turn sampler (NUTS) [Neal:2011mrf, Hoffman:2011ukg, Betancourt:2017ebh] that is implemented in the numpyro package [Phan:2019elc]. NUTS extends HMC to generate fairly uncorrelated samples by identifying a U-Turn in the trajectories of the HMC sampling process. No burn-in removal is required for this sampler as this accounted for by the so-called warm-up phase. During this phase (for which we choose a length of 2000 samples), one tunes the step-size and mass matrices of the integrator. We then draw 1500 posterior samples of our target parameters with three chains.
@@ 3 Simulated data
[…]
@@ 4.1.1 Validating the reconstruction (redshift + polar angle)
We start with marginalizing over the absolute magnitude coordinate [^6] We reconstruct the number counts up to an absolute magnitude of $M_{\rm thresh}=-19$. We determine in Sec. 4.4 the absolute magnitude above which the reconstruction can no longer be trusted. , and fix the azimuthal angle to 0 and $\pi$ in order to obtain the entire xzx-z plane, shown in Fig. 7. The panels are, from left to right, top to bottom: (1) The true galaxy number count, galaxies binned by their true redshift and polar angle, (2) the observed galaxy number count, binned by their measured redshift and polar angle, (3) the reconstructed galaxy Poisson rate in true redshift and polar angle, (4-6) the galaxy densities (same as panel (1-3) but dividing out the comoving volume), (7) the relative error of the galaxy Poisson rate, i.e.
$\text{Panel (7):}\quad\sigma\coloneqq\frac{{\rm std}(\mu(z,\varphi_{\perp};\Lambda))}{{\rm median}(\mu(z,\varphi_{\perp};\Lambda))}\,$, (4.1)
@@
$\text{Panel (8):}\quad\Delta\coloneqq n_{{\rm g},I}-\mu(z,\varphi_{\perp};\Lambda)\,.$ (4.2)
@@
$\text{Panel (9):}\quad\Delta_{\rm std}\coloneqq\frac{\Delta}{{\rm std}(\mu(z,\varphi_{\perp};\Lambda))}\,.$ (4.3)
@@ 4.1.4 The reconstructed line-of-sight redshift prior
To understand the implications of the reconstruction, we plot the true galaxy number count and inferred galaxy Poisson rate for a given HEALPixel, summing over the absolute magnitude counts. We focus on two different cases: (i) a pixel with a large number of observed galaxies, i.e. with a large apparent magnitude threshold, and (ii) a pixel in the “milky-way region”, i.e. with a small apparent magnitude threshold. The two cases are shown in the top and bottom panels of Fig. 10, respectively. There are two differences between the true and observed number counts: while they agree for low redshifts, the ratio of observed to true galaxy number counts drops for larger redshift due to fainter galaxies being missed. The second difference is that the observed number counts lack the small-scale fluctuations of the true number counts—this is due to the redshift uncertainties smearing out any fluctuations smaller than the standard deviation of the redshift distribution.
@@
Figure 11: Galaxy number counts in apparent magnitude and redshift for a selected pixel (top) and summed over the full sky (bottom). The columns from left to right are: the true number counts, the observed number count, the reconstructed Poisson rate, the relative difference, i.e. Eq. (4.4). The cut in the observed number counts at $m=22$ is due to the likelihood being evaluated on a grid up to this value – very few galaxies are observed above these apparent magnitudes. Note that the upper-left dark triangle is due to us excluding all magnitudes above 19-19, while the lower-right section is dark because no galaxies above a certain brightness exist. During inference, we do not impose any upper threshold on the absolute magnitude, this is only applied in post-processing.
@@ 4.1.5 The inferred absolute magnitude distribution
Let us first focus on the true catalog (first column). We see the apparent magnitude evolution (the approximately square root trend), reminiscent of Eq. (2.8) and Fig. 5. As galaxies are located at larger redshifts, they appear fainter (i.e. at larger $m$). We also notice an overall increase in the number counts with redshift, which is due to the larger associated comoving volume for voxels at larger redshifts. The second column shows the observed galaxy number counts, which are smaller than the true number counts due to the lower detection probability at larger apparent magnitudes. We have marked the apparent magnitude at which only $50\%$ of the galaxies are detected as a red dashed horizontal line. Above this line fewer galaxies are detected – and at an apparent magnitude of $m=22$ the observed catalog is essentially empty. Note that we have applied an absolute magnitude cut of $M_{\rm thr}=-19$, which is reflected in a dark region in the top left region of each plot. We decide to eliminate these galaxies (in post-processing) since they have low GW host probability and hence also contribute little to the redshift measurement for GWs. We will explore the impact of this cut in more detail in Sec. 4.4.
@@
Fig. 12 shows the absolute magnitude distribution of all (detected) galaxies in light blue (green), and the reconstructed galaxy Poisson rate in dark blue. The reconstruction recovers the full distribution faithfully, up to faint galaxies ($M_{h}\sim-18$), where $M_{h}$ is defined in Eq. (A.1). Since we reconstruct the galaxy rate only up to magnitudes above this value, the deviation between the true distribution and our reconstruction does not impact the final estimate of the number counts, as discussed in Sec. 4.4. The faithful recovery of the true magnitude distribution also justifies the generic galaxy magnitude modeling through a correlated Gaussian field that was introduced in App. A.1. It is universal enough to capture the non-trivial galaxy magnitude distribution of [Bertone:2007sj].
@@ 4.3 Impact of redshift uncertainties
For the simulation of the observed galaxy number counts, we assumed a Gaussian distribution for the redshift uncertainties, with an evolving standard deviation of $\sigma_{z}=0.01\times(1+z)$. This is a reasonable assumption for upcoming spectroscopic surveys [DESI:2013agm, DESI:2016fyo, LSSTScience:2009jmu], but it is important to understand how the results depend on the modeled redshift uncertainty as we assume it is constant. To this end, we repeat the analysis of Sec. 4.1 for different values of $\sigma_{z}$ (fixing $H_{0}=67\,\,{\rm km\,s^{-1}\,Mpc^{-1}}$). We assume (i) $\sigma_{z}=0.005$, (ii) $\sigma_{z}=0.02$ and (iii) $\sigma_{z}=0.07$, using the same simulated redshift uncertainties, i.e. $\sigma_{z}=0.01\times(1+z)$ in all cases. In Fig. 15, we show the deviations $\Delta_{\rm std}$ (cf. Eq. (4.3)) for each of these three analyses. As explained in a previous footnote, the $\Delta_{\rm std}$ distribution should follow approximately a zero-mean, unit-variance normal distribution. In the case of $\sigma_{z}=0.005$ (violet), the variance of the deviations is too large, indicating that the galaxy rate uncertainties are underestimated, as expected: the redshift uncertainties are too small when compared to the simulated uncertainties. In the case of $\sigma_{z}=0.02$ (blue), the deviations are significantly closer to the zero-mean unit-variance Gaussian distribution. Finally, in the case of $\sigma_{z}=0.07$ (yellow), the deviations show a variance that is too small – the uncertainties are overestimated.
We thus conclude that in order to be conservative, one should either (i) account for the redshift-dependent uncertainties or, (ii) use a larger redshift uncertainty overall that the effectively accounts for the redshift-evolving uncertainties. While the first option is the most accurate, it is also the most computationally expensive as the process to compute $\mu(\hat{z},\varphi_{\perp},m;\Lambda)$ will no longer be a convolution with a Gaussian kernel (cf. Eq. (2.11)). Depending on the rate of change of the width of the convolution with redshift, the computational cost can increase by an order of magnitude or more. Let us note here that the proposed framework is also suitable to analyze catalogs with both photometric and spectroscopic redshifts but requires a modified likelihood that accounts for the different redshift uncertainties. This analysis would also entail inferring two different selection effects, one for each measurement type.